Rework problem 21 from section 2.4 of your text, involving the assignment of tasks to student union board members. Assume that there are 11 board members: 7 females, and 4 males including Tom. There are 2 tasks to be assigned randomly, including that of reserving a room for meetings. There is at most one task per person. (1) Find the probability that Tom is given a task. equation editorEquation Editor (2) Find the probability that Tom is given the task of reserving a room for meetings.
Answer #1

task 1 : reserve room

no. of ways = (choose 1 person for task 1 )* (choose 1 person out of remaining for task 2 )

= 11 * 10 = 110

1.

no. of ways such that tom is assigned a task

= (tom will do task 1 )* (choose 1 person out of remaining for task 2 ) + (tom will do task 2 )* (choose 1 person out of remaining for task 1 )

= 1*10 + 1*10 = 20

P(tom is assigned task)

= 20/110

= 2/11

= 0.1818

2.

P(tom given task of reserving room (task 1))

= ( (tom will do task 1 )* (choose 1 person out of remaining for task 2 ) ) / 110

= 1*10 / 110

= 0.0909

(please UPVOTE)